# Metastability induced by non-reciprocal adaptive couplings in Kuramoto models

### v1

@@ III.1 Stationary states and choice of parameters
To visualize the dependence of collective states on the adaption rates $(\varepsilon_{1},\varepsilon_{2})$, we perform numerical simulations of the coupled system (1)-(3) with $N=10$ and $N=20$ oscillators over multiple realizations.
The simulations are carried out in the noise-free case with identical intrinsic frequencies $\omega_{i}=0$ for all oscillators $i=1,\dots,N$. To track the long-time dynamics, the order parameters, $R$ and $R_{2}$, are computed and averaged over the last sufficiently long time window. Each parameter pair $(\varepsilon_{1},\varepsilon_{2})$ is simulated for $5$ independent realizations, and the resulting mean order parameters $\langle R\rangle$ and $\langle R_{2}\rangle$ are used to classify the collective state.
The resulting categorical heatmap is shown in Fig. (1). Each pixel corresponds to a $(\varepsilon_{1},\varepsilon_{2})$ pair, color-coded by the observed collective state. Since $\varepsilon_{2}>0$ corresponds to the anti-Hebbian coupling, it tends to disrupt phase coherence, and sufficiently large values of $\varepsilon_{2}$ generally lead to incoherent dynamics. In contrast, when $\varepsilon_{1}\gg\varepsilon_{2}>0$, the upper-triangular Hebbian interactions dominate, promoting the formation of two anti-phase clusters for most parameter realizations.
To examine the dynamics in detail, we simulate the system (1)-(3) with $N=10$ oscillators, noise intensity $\sigma=0$ and intrinsic frequencies $\omega_{i}=0$ $\forall i$. We fix the coupling parameters at $\varepsilon_{1}=0.01$ and $\varepsilon_{2}=0.0001$, a regime where two anti-phase clusters are expected from Fig. (1) (a). While one might anticipate that $R_{2}$ would converge to one after an initial transient, we observe metastable behavior: $R_{2}$ initially approaches $1$, then occasionally deviates from this value, only to return to $R_{2}\approx 1$ after a few iterations. We have seen the same recurrent deviations of $R_{2}$ as shown in Fig. (2) (a) also over a total integration time of up to $9\times 10^{7}$ steps.
@@
Figure 3: Snapshots at $t=2000$ of the phases(a),the order parametres $R$ and $R_{2}$ (b) and the cluster sizes (c) before a switching event for $N=20$.
Figure 4: Same as Fig. 3, but snapshot at $t=4000$ after a switching event. Note that the cluster sizes have changed to a different partition before and after the switching of oscillators. In both panels (a) and (b), two anti-phase clusters are visible, each perfectly synchronized.
@@
It is instructive to plot the evolution of the asymmetric coupling matrix $k_{ij}$ in Fig. (2) (b-c). In panel (b), we observe the time evolution of the upper part $k_{ij}$ for $i<j$. As the system evolves, some of the couplings settle temporarily at +1+1, while others stabilize temporarily at 1-1. However, when $R_{2}$ deviates from $1$, there are switches in some of the $k_{ij}$ values from +1+1 to 1-1 or vice versa, indicating a temporary departure from the anti-phase synchronized two-cluster state. At the same time, some of the lower couplings $k_{ij}$ for $i>j$ also undergo slope changes, as shown in panel (c), but evolving on a much slower time scale. Panel (d) displays the evolution of the lower couplings on a long time scale, where it is seen more clearly that not all couplings are affected by a switch of $R_{2}$ and one may wonder whether some periodic structure is seen on long time scales.
@@
Metastability is the essential novel feature that we observe in these combinations of non-reciprocal couplings with respect to their sign and time scales. It is absent in the symmetric case [11]. Even, when $\varepsilon_{1}\gg\varepsilon_{2}$ with $\varepsilon_{1}=0.01$, $\varepsilon_{2}=0.0001$, but $\beta_{1}=\beta_{2}$, the system does not exhibit metastability. In this case, when $\beta_{1}=\beta_{2}=-\frac{\pi}{2}$, two stable anti-phase clusters emerge, while for $\beta_{1}=\beta_{2}=+\frac{\pi}{2}$, oscillations are only incoherent. Metastability, on the other hand, is observed in our system when there is sufficient asymmetry in the adaptive coupling dynamics, specifically when $\varepsilon_{1}\gg\varepsilon_{2}>0$, as shown in the orange-colored region of Fig. (1) as well as $\beta_{1}\neq\beta_{2}$. The slow dynamics adds as a perturbation on the fast dynamics; for the fast dynamics alone we would see a dynamic realization of Harari’s theorem [33]. According to that, the set of all points a balanced signed graph can be partitioned into two disjoint sets such that each positive line joins two points of the same subset and each negative line joins two points from different subsets. In our case, positive lines are realized as upper couplings within the same cluster and negative lines as upper couplings between the other cluster. The two anti-phase synchronized clusters are then perturbed by the slow dynamics which is manifest in some oscillators changing their cluster affiliation from time to time.
@@
Numerically we confirm that occasionally some oscillators switch groups, causing $R_{2}$ to deviate from 1. After a few steps, these oscillators return to a possibly slightly different anti-phase cluster configuration, and $R_{2}$ revisits the value of 1. The number of oscillators in each cluster may change with each deviation of $R_{2}$ from 1. Here, we include four videos [34] for different system sizes, $N=20$ and $N=200$, demonstrating these features, in particular how the group sizes may change with each spike in $R_{2}$. In Figs. 3 and  4 we include snapshots of these movies before and after a spike in $R_{2}$. What causes the changes in the slope of fast and slowly evolving couplings and which couplings are affected? Let us consider the simplest case, where a single oscillator $l$ changes its cluster affiliation. What happens to the couplings $k_{lk}$ in which this oscillator is involved? From the evolution of the coupling strengths $k_{ij}$ given in Eq. (3) it is seen that two oscillators with indices $l$ and $k$, chosen from the $N$ oscillators, which initially belong to the same cluster at time $t=t_{1}$ in an anti-phase cluster configuration with $R_{2}=1$ evolve as
@@
Numerically and visible in the attached movies we observe that depending on the system size, for small sizes $N=10$ or $N=20$, it is one or a few oscillators that change their cluster affiliation at the spikes of $R_{2}$, for larger sizes these are more. What is typical in the movies is the splitting into two clusters of similar size as well as the fact that only a few oscillators change their affiliation. This reminds to “weak ties” in social systems, where a few members of a community are weakly bound and willing to switch their community (of shared opinions, for example), while the majority is strongly bound, and the weak ties are of beneficial impact on the system, as analyzed by Gronevetter [35]. The reason why oscillators change their cluster affiliation is determined by the attractor landscape and the saddles generated by the dynamics of Eqs. 1-3 that we analyze in section III.4. In brain dynamics this would correspond to oscillatory units which easily switch the group with which they synchronize.
@@ III.2 Features of randomness
The time evolution of the order parameters looks like generated by a stochastic source in spite of the fully deterministic dynamics of Eqs. 13. The events when one or more oscillators change their cluster affiliation look randomly distributed. The only source of randomness is the quenched distribution of initial phases and couplings which determine the initial distance of the phases from the saddle equilibria and therefore their trajectories through attractor space. Actually, on an intermediate time scale of the order of the interval between two switching events, the distance from the next saddle is essential to predict which oscillator escapes first in a next switching event. Assuming that there is a switch at time $t$ of one or a few oscillators and that we know the instantaneous phase velocities $\dot{\theta}_{i}$ at an earlier time $t_{snap}=t-\tau_{win}$ from a snapshot at that time, we approximate the estimated time of escape (ETE) for each oscillator $i$ as
@@
Note that the snapshot of velocities $\dot{\theta}_{i}(t_{\text{snap}})$ depends on the upper and lower couplings which coevolve with the phases. However, our previous figures for $k_{ij}$ show that within two switching events upper couplings are almost constant, as they converge fast to their equilibrium values, and lower couplings are so slow that their change may be neglected. Moreover, the distance is assumed to increase linearly with time. Nevertheless this approximation works quite well to predict the dynamics after a switch including an identification which oscillator will switch first. Before we discuss the concrete results, some remarks are in order about the choice of $\tau_{{\text{win}}}$, $\tau_{{\text{win}}}=200$ turns out to be a good choice, given the other parameters. If $\tau_{{\text{win}}}$ is chosen too large, the co-evolution of couplings can no longer be neglected, oscillators whose ETE exceeds a cutoff value $\eta_{\max}=500$ steps are excluded from consideration. If $\tau_{{\text{win}}}$ is too short, there is no prediction. If more than one oscillator switches its cluster affiliation, we group individual switches that happen very close together into a single collective event. If several oscillators overcome the distance of $\pi$ within $50$ simulation steps of each other, they are evaluated as a single event. For entire groups changing their affiliation, we predict the membership of the switching cluster $\mathcal{P}$ by including any oscillator $j$ whose ETE lies within a temporal margin $\Delta\eta$ of the “leading oscillator” $i$: $\mathcal{P}=\{j\mid\eta_{j}\leq\eta_{i^{*}}+\Delta\eta\}$. This margin accounts for the finite temporal spread of collective switching events and is set to $\Delta\eta=150$ steps in our simulations, based on the typical duration of observed group switching.
@@
What cannot be predicted is the new partition into two clusters after the switching event and the dynamics on a long-time scale. One might be interested in the probability that a given set of oscillator phases find the same couplings (actually coupling conditions for their own evolution) after some possibly long time, as the evolution -locally in time- looks random. The question for finding repeatedly the same conditions is the question for long-time periodicity. Such long periods in the Kuramoto order parameter from short period oscillations have been observed in [36]. For larger networks we cannot answer this question, but for a system of $N=4$ oscillators, the six upper couplings and the six lower couplings synchronize into one or two clusters, as well as the phases, for $N=4$ as for $N=3$ the system has only one saddle point and the evolution of all couplings is strictly periodic, see Fig. 6.
In summary, on intermediate time scales it is the momentary distance from $\pi$ of the phase position and the phase velocity at that instant of time, which allow a prediction of which oscillator (group of oscillators) will escape next, knowing that there is an escape event. As soon as this distance is overcome, we can be sure to be in the attraction regime of the other cluster. It is the random initial conditions which determine the initial phase position and lead to an ensemble of time evolutions unless identical initial conditions are chosen via the same seed. Given a set of fixed initial conditions, the waiting time distribution between two switches of the order parameter $R_{2}$ looks random on longer time scales, unless it happens to become periodic. We address this irregularity to the complex attractor landscape that allows the escape from saddles and determines the entrance to the attraction regimes of new saddles.
@@ III.3 Choice of initial conditions and natural frequencies
If we choose uniformly distributed initial configurations for upper and lower couplings very close to 1, this choice results in some delay of the onset of metastable dynamics, where for the upper couplings it is the fully synchronized state that is stable until metastability sets in, see Fig. 7.
@@ III.4 Metastability as a function of the system size
For $N=200$ the distinction between transient stable configurations interrupted by short excursions to other clusters becomes less pronounced. The movie of [34] for $N=200$ shows the evolution of phases, snapshots are given in Figs. 9 and 10 before and after a spike in $R_{2}$.
Figure 9: Same as Fig. 3, but for $N=200$. Note that two anti-phase clusters are less perfectly synchronized than for $N=20$. The partition into two clusters is roughly half-half.
@@
We still see remnants of an organization into two anti-phase synchronized clusters, where on short time scales always some oscillators escape and return to their cluster, more frequent are the events where the escape succeeds over a distance $\pi$ in the phases.
Figure 10: Same as Fig. 4, but for $N=200$. The size of the two clusters of the partition varies only slightly before and after the switch, but as compared to $N=20$, more and more oscillators are involved in changing their cluster affiliation .
At intermediate times also configurations with several coexisting clusters are forming, leading to stronger spikes in $R_{2}$.
@@
To understand more quantitatively the increasing instability with the system size, we summarize in the following the stability analysis which is detailed in the Appendix. For the fully connected Kuramoto model with $\omega_{i}=0$ and $\sigma=0$ we start from a two-cluster anti-phase synchronized solution with clusters A and B of size $N_{A}$ and $N_{B}$. Next we perform a stability analysis around the two-cluster solution. The Jacobian takes a block form with four blocks $J_{\theta\theta}$, $J_{\theta k}$, $J_{k\theta}$, and $J_{kk}$. For the phase-coupling block $J_{\theta k}$ it is easy to show that the matrix vanishes. Therefore the determinant of the entire Jacobian factorizes into those of the diagonal entities $\det{J_{\theta\theta}}\det{J_{kk}}$, and the eigenvalues are the union of those of $J_{\theta\theta}$ and $J_{kk}$. The eigenvalues of $J_{kk}$ turn out to be all negative. Thus the problem is reduced to determine the eigenvalues of $J_{\theta\theta}$. At the two-cluster anti-phase equilibrium the couplings attain their equilibrium values $\dot{k}_{ij}=0$, which leads to vanishing row sums and one zero eigenvalue, corresponding to the global phase shift of the two-cluster solution. Moreover, it can be shown that $\chi J_{\theta\theta}\chi=-J_{\theta\theta}$ with a flip matrix $\chi$. Therefore the eigenvalues of $J_{\theta\theta}$ occur in symmetric pairs with m1m-1 positive values $\lambda_{k}=1-2k/N,k=1,2,...,m-1$ and m1m-1 negative values $\lambda_{-k}=-(1-2k/N),k=1,2,...,m-1$ for $N=2m$ plus $2$ zero eigenvalues; for $N=2m+1$ we have $m$ positive and $m$ negative values and one zero eigenvalue, $\lambda_{k},\lambda_{-k}$ as before with k=1,2,mk=1,2,...m. This proves that with increasing system size $N$, we have an increasing number of unstable directions for the two anti-phase synchronized cluster solution.

### v2

@@ III.1 Stationary states and choice of parameters
To visualize the dependence of collective states on the adaptation rates $(\varepsilon_{1},\varepsilon_{2})$, we perform numerical simulations of the coupled system (1)-(2) with $N=10$ and $N=20$ oscillators over multiple realizations.
The simulations are carried out in the noise-free case with identical intrinsic frequencies $\omega_{i}=0$ for all oscillators $i=1,\dots,N$. To track the long-time dynamics, the order parameters, $R$ and $R_{2}$, are computed and averaged over the last sufficiently long time window. Each parameter pair $(\varepsilon_{1},\varepsilon_{2})$ is simulated for $200$ independent realizations, and the resulting mean order parameters $\langle R\rangle$ and $\langle R_{2}\rangle$ are used to classify the collective state. Specifically for the construction of these phase diagrams, where covering the parameter space with high resolution requires massive computational throughput ($200$ realizations per grid point), we used a shorter simulation time of $4\times 10^{4}$ steps. The order parameters were averaged over the final $1000$ steps after the system had relaxed into its attractor. We verified that this duration is sufficient for the collective state (two anti-phase clusters vs. incoherent) to stabilize.
The resulting categorical heatmap is shown in Fig. (1). Each pixel corresponds to an $(\varepsilon_{1},\varepsilon_{2})$ pair, color-coded by the observed collective state. Since $\varepsilon_{2}>0$ corresponds to the anti-Hebbian coupling, it tends to disrupt phase coherence, and sufficiently large values of $\varepsilon_{2}$ generally lead to incoherent dynamics. In contrast, when $\varepsilon_{1}\gg\varepsilon_{2}>0$, the upper-triangular Hebbian interactions dominate, promoting the formation of two anti-phase clusters for most parameter realizations. For $\varepsilon$-values in the vicinity of the red dashed line it depends on the random initial conditions in which state the system ends up, therefore the boundary is a bit rough.
To examine the dynamics in detail, we simulate the system (1)-(2) with $N=10$ oscillators, noise intensity $\sigma=0$ and intrinsic frequencies $\omega_{i}=0$ $\forall i$. We fix the coupling parameters at $\varepsilon_{1}=0.01$ and $\varepsilon_{2}=0.0001$, a regime where two anti-phase clusters are expected from Fig. (1) (a). While one might anticipate that $R_{2}$ would converge to one after an initial transient, we observe metastable behavior: $R_{2}$ initially approaches $1$, then occasionally deviates from this value, only to return to $R_{2}\approx 1$ after a few iterations. We observe the same recurrent deviations of $R_{2}$ as shown in Fig. (2) (a) also over a total integration time of up to $9\times 10^{7}$ steps.
@@
It is instructive to plot the evolution of the asymmetric coupling matrix $k_{ij}$ in Fig. (2) (b-c). In panel (b), we observe the time evolution of the upper part $k_{ij}$, that is, for $i<j$. As the system evolves, some of the couplings settle temporarily at +1+1, while others stabilize temporarily at 1-1. However, when $R_{2}$ deviates from $1$, there are switches in some of the $k_{ij}$ values from +1+1 to 1-1 or vice versa, indicating a temporary departure from the anti-phase synchronized two-cluster state. At the same time, some of the lower couplings $k_{ij}$, that is, for $i>j$ also undergo slope changes, as shown in panel (c), but evolving on a much slower time scale. Note that the transient dynamics of the order parameters $R$ and $R_{2}$ appear slightly faster even than those of the upper couplings $k_{ij}$ (Fig. 2 (a-b)). This is due to the timescale separation explicitly built into the model: the phases $\theta_{i}$ evolve on a timescale of $O(1)$ (Eq.  (1)), whereas the upper part of the couplings evolve with rate $\varepsilon_{1}=0.01$. Consequently, the oscillators can phase-lock and stabilize $R$ and $R_{2}$ relatively quickly, while the coupling weights require some integration time to saturate to their equilibrium values. Panel (d) displays the evolution of the lower couplings on a long time scale, where it becomes more clearly visible that not all couplings are affected by a switch of $R_{2}$. Here one may wonder whether some periodic structure is seen on long time scales.
@@
Metastability as an essential novel feature. Metastability is the essential novel feature that we observe in these combinations of non-reciprocal couplings with respect to their sign and time scales. It is absent in the reciprocal case [20]. Even, when $\varepsilon_{1}\gg\varepsilon_{2}$ with $\varepsilon_{1}=0.01$, $\varepsilon_{2}=0.0001$, but $\beta_{1}=\beta_{2}$, the system does not exhibit metastability. In this case, when $\beta_{1}=\beta_{2}=-\frac{\pi}{2}$, two stable anti-phase clusters emerge, while for $\beta_{1}=\beta_{2}=+\frac{\pi}{2}$, oscillations are only incoherent. Metastability, on the other hand, is observed in our system when there is sufficient asymmetry in the adaptive coupling dynamics, specifically when $\varepsilon_{1}\gg\varepsilon_{2}>0$, as shown in the orange-colored region of Fig. (1) as well as $\beta_{1}\neq\beta_{2}$. The slow dynamics acts as a perturbation on the fast dynamics; for the fast dynamics alone we would see a kind of dynamic realization of Harary’s theorem [47]. According to that, the set of all points of a balanced signed graph can be partitioned into two disjoint sets such that each positive line joins two points of the same subset and each negative line joins two points from different subsets, the lines are undirected. In our case, we have directed links with different dynamics assigned to both directions; but as long as the lower couplings are neglected which act in opposite link direction to the upper ones, what dynamically matters is which dynamics is assigned to the lines: these are attractive upper couplings, corresponding to positive lines, and realized as zero phase difference within the same cluster, as well as repulsive upper couplings, corresponding to negative lines, and realized as phase difference of $\pi$ between members of different clusters. The two anti-phase synchronized clusters are then slightly perturbed by the slow dynamics of lower couplings which is manifest in some oscillators changing their cluster affiliation from time to time.
@@
Switching the cluster affiliation. Numerically we confirm that occasionally some oscillators switch groups, causing $R_{2}$ to deviate from 1. After a few steps, these oscillators return to a possibly slightly different anti-phase cluster configuration, and $R_{2}$ revisits the value of 1. The number of oscillators in each cluster may change with each deviation of $R_{2}$ from 1. Here, we include four videos [48] for different system sizes, $N=20$ and $N=200$, demonstrating these features, in particular how the group sizes may change with each spike in $R_{2}$. In Fig. (3), we include snapshots of these movies before and after a spike in $R_{2}$.
What causes the changes in the slope of fast and slowly evolving couplings and which couplings are affected? Let us consider the simplest case, where a single oscillator $l$ changes its cluster affiliation. What happens to the couplings $k_{lk}$ in which this oscillator is involved? From the evolution of the coupling strengths $k_{ij}$ given in Eq. (2) it is seen that two oscillators with indices $l$ and $k$, chosen from the $N$ oscillators, which initially belong to the same cluster at time $t=t_{1}$ in an anti-phase cluster configuration with $R_{2}=1$ evolve as
@@
Numerically and visible in the attached movies we observe that depending on the system size, for small sizes $N=10$ or $N=20$, it is a single or a few oscillators that change their cluster affiliation at the spikes of $R_{2}$, for larger sizes these are more. What is typical in the movies is the splitting into two clusters of similar size as well as the fact that only a few oscillators change their affiliation. This reminds is reminiscent to “weak ties” in social systems, where a few members of a community are weakly bound and willing to switch their community (of shared opinions, for example), while the majority is strongly bound, and the weak ties are of beneficial impact on the system, as analyzed by Granovetter [49]. The reason why oscillators change their cluster affiliation is determined by the attractor landscape and the saddles generated by the dynamics of Eqs. (1)-(2) that we analyze in section (III.4). In brain dynamics this would correspond to oscillatory units which easily switch the group with which they synchronize, where a single unit may represent a population of neurons.
@@ III.2 Features of randomness
The time evolution of the order parameters appears to be generated by a stochastic source in spite of the fully deterministic dynamics of Eqs. (1)- (2). The events when one or more oscillators change their cluster affiliation look randomly distributed. The only source of randomness is the quenched distribution of initial phases and couplings which determine the initial distance of the phases from the saddle equilibria and therefore their trajectories through attractor space. Actually, on an intermediate time scale of the order of the interval between two switching events, the distance from the next saddle is essential to predict which oscillator escapes first in a next switching event. Assuming that there is a switch at time $t$ of one or a few oscillators and that we know the instantaneous phase velocities $\dot{\theta}_{i}$ at an earlier time $t_{snap}=t-\tau_{win}$ from a snapshot at that time, we approximate the estimated time of escape (ETE) for each oscillator $i$ as
@@
Note that the snapshot of velocities $\dot{\theta}_{i}(t_{\text{snap}})$ depends on the upper and lower couplings which coevolve with the phases. However, our previous figures for $k_{ij}$ show that within two switching events upper couplings are almost constant, as they converge fast to their equilibrium values, and lower couplings are so slow that their change may be neglected. Moreover, the distance is assumed to increase linearly with time. Nevertheless this approximation works quite well to predict the dynamics after a switch including an identification which oscillator will switch first.
Before we discuss the concrete results, some remarks are in order about the choice of $\tau_{{\text{win}}}$, $\tau_{{\text{win}}}=200$ turns out to be a good choice, given the other parameters. If $\tau_{{\text{win}}}$ is chosen too large, the co-evolution of couplings can no longer be neglected, oscillators whose ETE exceeds a cutoff value $\eta_{\max}=500$ steps are excluded from consideration. If $\tau_{{\text{win}}}$ is too short, there is no prediction. If more than one oscillator switches its cluster affiliation, we group individual switches that happen very close together into a single collective event. If several oscillators overcome the distance of $\pi$ within $50$ simulation steps of each other, they are evaluated as a single event. For entire groups changing their affiliation, we predict the membership of the switching cluster $\mathcal{P}$ by including any oscillator $j$ whose ETE lies within a temporal margin $\Delta\eta$ of the “leading oscillator” $i$: $\mathcal{P}=\{j\mid\eta_{j}\leq\eta_{i^{*}}+\Delta\eta\}$. This margin accounts for the finite temporal spread of collective switching events and is set to $\Delta\eta=150$ steps in our simulations, based on the typical duration of observed group switching.
@@
What cannot be predicted is the new partition into two clusters after the switching event and the dynamics on a long-time scale. One might be interested in the probability that a given set of oscillator phases finds the same couplings (actually coupling conditions for their own evolution) after some possibly long time, as the evolution -locally in time- looks random. The question for finding repeatedly the same conditions is the question for long-time periodicity. Such long periods in the Kuramoto order parameter from short period oscillations have been observed in [50]. For larger networks we cannot answer this question, but for a system of $N=4$ oscillators, the six upper couplings and the six lower couplings synchronize into one or two clusters, as well as the phases, both for $N=4$ and $N=3$ the system has only one saddle point and the evolution of all couplings is strictly periodic, see Fig. 5.
In summary, on intermediate time scales it is the momentary distance from $\pi$ of the phase position and the phase velocity at that instant of time, which allow a prediction of which oscillator (group of oscillators) will escape next, knowing that there is an escape event. As soon as this distance is overcome, we can be sure to be in the attraction regime of the other cluster. It is the random initial conditions which determine the initial phase position and lead to an ensemble of time evolutions unless identical initial conditions are chosen via the same seed. Given a set of fixed initial conditions, the waiting time distribution between two switches of the order parameter $R_{2}$ looks random on longer time scales, unless it happens to become periodic. We address this irregularity to the complex attractor landscape that allows the escape from saddles and determines the entrance to the attraction regimes of new saddles. The entrance location determines the dwell time in the vicinity of the new saddles.
Figure 8: Snapshots of the system state before and after a switching event for the larger system size $N=200$. (a–c) State at $t=820$ (before the switch). (a) Instantaneous phases on the unit circle; note that the two clusters are less perfectly synchronized than in the $N=20$ case. (b) Time evolution of order parameters. (c) Cluster sizes showing a partition of $115$ and $85$ oscillators. (d–f) State at $t=1500$ (after the switch). (d) Instantaneous phases. (e) Evolution of order parameters. (f) Cluster sizes have shifted to a partition of $109$ and $91$ oscillators. Compared to $N=20$, the fluctuations in cluster size are more frequent, involving multiple oscillators switching affiliation.
@@ III.3 Choice of initial conditions and natural frequencies
If we choose uniformly distributed initial configurations for upper and lower couplings very close to 1, this choice results in some delay of the onset of metastable dynamics, where for the upper couplings it is the fully synchronized state that remains stable until metastability sets in, see Fig. 6.
@@ III.4 Metastability as a function of the system size
For $N=200$ the distinction between temporarily stable configurations interrupted by short excursions to other clusters becomes less pronounced. The movie of [48] for $N=200$ shows the evolution of phases, snapshots are given in Fig. (8) before and after a spike in $R_{2}$.
@@
We still see remnants of an organization into two anti-phase synchronized clusters, where on short time scales always some oscillators escape and return to their cluster, more frequent are the events where the escape succeeds over a distance $\pi$ in the phases. At intermediate times also configurations with several coexisting clusters are forming, leading to stronger spikes in $R_{2}$.
@@
To understand more quantitatively the increasing instability with the system size, we summarize the stability analysis in the following . It is detailed in the Appendix. For the fully connected Kuramoto model with $\omega_{i}=0$ and $\sigma=0$ we start from a two-cluster anti-phase synchronized solution with clusters A and B of size $N_{A}$ and $N_{B}$. Next we perform a stability analysis around the two-cluster solution. The Jacobian takes a block form with four blocks $J_{\theta\theta}$, $J_{\theta k}$, $J_{k\theta}$, and $J_{kk}$. For the phase-coupling block $J_{\theta k}$ it is easy to show that the matrix vanishes. Therefore the determinant of the entire Jacobian factorizes into those of the diagonal entities $\det{J_{\theta\theta}}\det{J_{kk}}$, and the eigenvalues are the union of those of $J_{\theta\theta}$ and $J_{kk}$. The eigenvalues of $J_{kk}$ turn out to be all negative. Thus the problem is reduced to determine the eigenvalues of $J_{\theta\theta}$. At the two-cluster anti-phase equilibrium the couplings attain their equilibrium values $\dot{k}_{ij}=0$, which leads to vanishing row sums and one zero eigenvalue, corresponding to the global phase shift of the two-cluster solution. Moreover, it can be shown that $\chi J_{\theta\theta}\chi=-J_{\theta\theta}$ with a flip matrix $\chi$. Therefore the eigenvalues of $J_{\theta\theta}$ occur in symmetric pairs with m1m-1 positive values $\lambda_{k}=1-2k/N,k=1,2,...,m-1$ and m1m-1 negative values $\lambda_{-k}=-(1-2k/N),k=1,2,...,m-1$ for $N=2m$ plus $2$ zero eigenvalues; for $N=2m+1$ we have $m$ positive and $m$ negative values and one zero eigenvalue, $\lambda_{k},\lambda_{-k}$ as before with k=1,2,mk=1,2,...m. This proves that with increasing system size $N$, we have an increasing number of unstable directions for the two anti-phase synchronized cluster solution.
What does this scaling mean in view of possible neural network applications? As stated in the introduction, individual Kuramoto oscillators are often used to model entire brain regions such as cortical areas, each treated as one oscillatory unit. For small system sizes such as $N=10$ or $N=20$ this could mean that one observes a pronounced metastability of synchronization patterns between brain areas. For a long dwell time they are synchronized in one pattern of two anti-phase synchronized clusters of areas, until suddenly one or a few of these areas switch their cluster affiliation: after the switch, they synchronize with the cluster complementary to the one they were synchronized with before the switch. The dwell time in one synchronization pattern decreases with the number of coupled areas, the more areas, the shorter the dwell time. We leave it open which assignment to a single Kuramoto oscillator is most appropriate, ranging from a single neuron to global areas in the brain.